Constraints on Growth: Bulla Recording, Inc., wishes to maintain a growth rate of 12 percent per...

  1. Constraints on Growth: Bulla Recording, Inc., wishes to maintain a growth rate of 12 percent per year and a debt-equity ratio of .30. Profit margin is 5.9 percent, and the ratio of total assets to sales is constant at .85. Is this growth rate possible? To answer, determine what the dividend payout ratio must be. How do you interpret the result?
Answer #1

According to DuPont Identity:

ROE= Profit margin*Asset turnover*Financial Leverage

Asset turnover=revenue/Assets=1/0.85

Financial leverage = Assets/Equity:
Debt/Equity=0.30

Debt=0.30
Equity=1

Asset=1+0.30=1.30

Asset/Equity=1.30/1=1.30

ROE=5.90%*1/0.85*1.30=9.0235%

Now:

Sustainable growth rate = ROE*retention ratio
12%=9.0235*retention ratio
Retention ratio=1.33

This retention ratio is not possible hence a growth rate of more than ROE is not possible as retention ratio cannot be more than 1. In this case dividend payout ratio shall be 1-1.33=-0.33
Which is not possible.

This means that the company has to recall the previous dividends from investors which has never been done and the investors do not have the obligations to do so if the company recalls dividend.